OpenStudy (anonymous):

Can someone help to walk me through the steps of 2 problems that I don't understand. It was a part of a review test. log,2x+3,(125)=3 and How long it will take for $300 to double when invested at 6% annual interest compounded twice a year? Thanks :)

OpenStudy (anonymous):

Is this the first one? \[\log_{2x+3}125=3\]

OpenStudy (anonymous):

2.let x be the no. of years,then \[600=300(1+\frac{ 6 }{ 200 })^{2*x}\] find x